What are 4 consecutive integers that total 856?
Answer by Alan3354(69443)

What are 4 consecutive integers that total 856?
------------
856/4 = 214, the center
212 + 213 + 214 + 215 = 854
213 + 214 + 215 + 216 = 858
-----
No such numbers to total 856